a spaceship has a weight of 8.82 x 104 N on the surface of the Earth. the space ship is launched from earth and lands on a distant planet where it has a weight of 3.90 x 105 N. what is the acceleration due to gravity on this planet?


1
Expert's answer
2021-04-23T11:11:10-0400
"g'=g\frac{W'}{W}\\g'=9.8\frac{39}{8.82}=43.3\frac{m}{s^2}"